OpenStudy (anonymous):

A set of sample data has a mean of 62 and a standard deviation of 7. Approximately what percentage of the data is between 41 and 83?

OpenStudy (zarkon):

Do we know anything about the underlying distribution?

OpenStudy (anonymous):

no. but one of the answers is not enough info provided

OpenStudy (zarkon):

If we knew that the distribution was bell shaped we could answer the question using the empirical rule.

OpenStudy (anonymous):

i think we could assume that it is

OpenStudy (zarkon):

99.7%

OpenStudy (anonymous):

can you tell me what you did?

OpenStudy (zarkon):

41 and 83 are three standard deviation from the mean...empirical rule gives us that we have approx 99.7% of data

OpenStudy (zarkon):

41=62-3*7 83=62+3*7